OpenStudy (anonymous):

√4:(x^(16)) use rational exponents to simplify.

OpenStudy (anonymous):

\[2x^8\]

OpenStudy (anonymous):

\[\sqrt{4x^16}=(4x^{16})^\frac{1}{2}=(2^2x^{16})^\frac{1}{2}\]\]

OpenStudy (anonymous):

multiply the exponents to get \[2x^8\]

OpenStudy (anonymous):

4x16−−−−√=(4x16)12=(22x16)12